# Optimizing EPR pulses for broadband excitation and refocusing

### v1

@@ VI Conclusion
In this paper, we have shown that it is possible to obtain about a $10\%$ improvement of the Hahn echo amplitude over an optimal, generalized hyperbolic secant KBB sequence by optimizing nonlinear amplitude-limited Fourier series, discrete time series, or neural network parameterized pulses. With these parameterized pulses, we find a 1:1 pulse length ratio is favorable in the presence of power constraints because it allows more energy to be allotted to the $\pi$ pulse while the pulse shaping is still able to maintain the refocusing. Interestingly, the optimization landscape for this type of problem has many equivalent minima. We find no marked differences among the three parameterizations. Furthermore, we have demonstrated that nonlinear effects due to amplifier compression and resonator transfer can be included in the optimization workflow, allowing for the usage of the full power of an amplifier, including its nonlinear region. These results demonstrate a pathway towards optimal broadband spectral acquisition with constrained power.

### v2

@@ VI Conclusion
In this paper, we have shown that it is possible to obtain about a $10\%$ improvement of the Hahn echo amplitude over an optimal, generalized hyperbolic secant KBB sequence by optimizing nonlinear amplitude-limited Fourier series, discrete time series, or neural network parameterized pulses. With these parameterized pulses, we find a 1:1 pulse length ratio is favorable in the presence of power constraints because it allows more energy to be allotted to the $\pi$ pulse while the pulse shaping is still able to maintain the refocusing. Interestingly, the optimization landscape for this type of problem has many equivalent maxima, all of which involve the $\pi/2$ and $\pi$ pulses cooperatively working together to compensate phase dispersions in each other. We find no marked differences among the three parameterizations. Furthermore, we have demonstrated that nonlinear effects due to amplifier compression and resonator transfer can be included in the optimization workflow, allowing for the usage of the full power of an amplifier, including its nonlinear region. These results demonstrate a pathway towards optimal broadband spectral acquisition with constrained power.